Perform the indicated operations.
Use three unit multipliers to convert 16,000 cubic centimeters to cubic meters.

Answer by josmiceli(19441)

(16000 cm3)(1 m/100 cm)(1m/100 cm)(1m/100 cm) = (16000/100^3 m3)
Note that 100^3 = 100*100*100
16000/100^3 m3 = .016 m3 answer